# Provably Efficient Reinforcement Learning with Multinomial Logit Function Approximation

### v2

@@ 3 Problem Setup
There exists $10<\kappa\leq\kappa^{*}<$ 1 such that for all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$ and $s^{\prime},s^{\prime\prime}\in\mathcal{S}_{h,s,a}$, it holds that $\inf_{\theta\in\Theta}p_{s,a}^{s^{\prime}}(\theta)p_{s,a}^{s^{\prime\prime}}(\theta)\geq\kappa$ and $p_{s,a}^{s^{\prime}}(\theta^{*})p_{s,a}^{s^{\prime\prime}}(\theta^{*})>\kappa^{*}$.
@@
Assumption 1 is similar to the standard assumption in generalized linear model literature (Filippi et al., 2010; Wang et al., 2021) and logistic bandits literature (Faury et al., 2020; Abeille et al., 2021; Zhang and Sugiyama, 2023) to guarantee the Hessian matrix is non-singular in Property 1.
4 Computationally Efficient Algorithm by Online Learning
In this part, we design an algorithm that achieves $\mathcal{O}(1)$ computational and storage costs per episode.
@@ 4 Computationally Efficient Algorithm by Online Learning
Since the transition parameter $\theta_{h}^{*}$ is unknown, we need to estimate it using the historical data. At episode $k$, we collect a trajectory $\{(s_{k,h},a_{k,h})\}_{h=1}^{H}$, then define the variable: $y_{k,h}\in\{0,1\}^{N_{k,h}}$ where $y_{k,h}^{s^{\prime}}=\operatorname{\mathds{1}}{(s_{k,h+1}=s^{\prime})}$ for $s^{\prime}\in\mathcal{S}_{k,h}\triangleq\mathcal{S}_{s_{k,h},a_{k,h}}$ and $N_{k,h}=\lvert\mathcal{S}_{k,h}\rvert$. We denote by $p_{k,h}^{s^{\prime}}(\theta)=p_{s_{k,h},a_{k,h}}^{s^{\prime}}(\theta)$. Then $y_{k,h}$ is a sample from the following multinomial distribution:
$y_{k,h}\sim\textnormal{multinomial}(1,[p_{k,h}^{s_{1}}(\theta^{*}),\ldots,p_{k,h}^{s_{N_{k,h}}}(\theta^{*})]),$
@@
where the parameter $1$ indicates that $y_{k,h}$ is a single-trial sample. Furthermore, we define the noise $\epsilon_{k,h}^{s^{\prime}}=y_{k,h}^{s^{\prime}}-p_{k,h}^{s^{\prime}}(\theta_{h}^{*})$. It is clear that $\epsilon_{k,h}\in[-1,1]^{N_{k,h}}$, $\mathbb{E}[\epsilon_{k,h}]=\mathbf{0}$ and $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\epsilon_{i,h}^{s^{\prime}}=0$.
Efficiency Concern.  Hwang and Oh (2023) made the first step by proposing an approach using the maximum likelihood estimation (MLE). Specifically, their estimator $\theta_{k,h}$ is defined as
@@
$\theta_{k,h}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}-y_{i,h}^{s^{\prime}}\log p_{i,h}^{s^{\prime}}(\theta)+\frac{\lambda}{2}\lVert\theta\rVert_{2}^{2}.$ (1)
Despite favorable theoretical guarantee, the computational and storage cost of this method is expensive. First, in episode $k$, the estimator $\theta_{k,h}$ in (1) is computed using all samples collected in previous episodes, resulting an $\mathcal{O}(k)$ storage cost. Second, to solve this optimization problem, the projected gradient descent (Boyd and Vandenberghe, 2004) are usually applied. However, as discussed in Faury et al. (2022), the optimization of the MLE problem typically require $\mathcal{O}(k\log(1/\epsilon))$ iterations to achieve an $\epsilon$-accurate solution, which is computationally expensive.
To address this issue, we introduce a novel algorithm named UCRL-MNL with Online Learning (UCRL-MNL-OL), which achieves both the same regret with $\mathcal{O}(1)$ computational and storage cost. At a high level, our algorithm can be divided into two phases: (i) efficient online estimation, and (ii) efficient optimistic value function construction. We introduce the details in the following.
Efficient Online Estimation.  Instead of using all historical data, our algorithm updates the estimator in an online manner. Specifically, inspired by the works (Hazan et al., 2014; Zhang et al., 2016) on logistic bandit, we find the negative log-likelihood function is exponentially concave, which motivates us to apply a variant of online Newton step (Zhang et al., 2016; Oh and Iyengar, 2021).
First, we define per-episodic loss function $f_{k,h}(\theta)$, gradient $g_{k,h}(\theta)$ and Hessian matrix $H_{k,h}(\theta)$ as
$f_{k,h}(\theta)$ $=-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}y_{k,h}^{s^{\prime}}\log p_{k,h}^{s^{\prime}}(\theta),\quad g_{k,h}(\theta)=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}},$
$H_{k,h}(\theta)$ $=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{s^{\prime\prime}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime\prime}})^{\top}.$ (2)
We introduce the following key property of the loss function $f_{k,h}(\theta)$ by a second-order Taylor expansion and mean value theorem, whose proof can be found in Lemma 5 in Appendix A.
Property 1.
(i). For any $\theta_{1},\theta_{2}\in\mathbb{R}^{d}$, $\exists~{}\bar{\theta}=\nu\theta_{1}+(1-\nu)\theta_{2}$ such that
$f_{k,h}(\theta_{2})=f_{k,h}(\theta_{1})+\langle g_{k,h}(\theta_{1}),\theta_{2}-\theta_{1}\rangle+\frac{1}{2}\lVert\theta_{2}-\theta_{1}\rVert_{H_{k,h}(\bar{\theta})}^{2}.$
(ii). For any $\theta\in\Theta$, it holds that
$H_{k,h}(\theta)\succeq\kappa\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}\triangleq\kappa W_{k,h}.$ (3)
Property 1 implies that $f_{k,h}(\theta)$ is exponentially concave, which enables us to apply the online Newton step to construct an efficient estimator. Specifically, we construct $\widehat{\theta}_{k+1,h}$ by solving the problem:
$\widehat{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\langle g_{k,h}(\widehat{\theta}_{k,h}),\theta-\widehat{\theta}_{k,h}\rangle+\frac{1}{2}\lVert\theta-\widehat{\theta}_{k,h}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2},$ (4)
where $\widehat{\Sigma}_{k+1,h}=\widehat{\Sigma}_{k,h}+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$ is the feature covariance matrix.
We show the online estimator $\widehat{\theta}_{k,h}$ in (4) enjoys computational and storage efficiency simultaneously. As the optimization problem exhibits a standard online mirror descent formulation, it can be solved with a projected gradient step with the following equivalent formulation by
$\theta^{\prime}_{k+1,h}=\widehat{\theta}_{k,h}-\widehat{\Sigma}_{k+1,h}^{-1}g_{k,h}(\widehat{\theta}_{k,h}),\quad\mbox{and}\quad\widehat{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\lVert\theta-\theta^{\prime}_{k+1,h}\rVert_{\widehat{\Sigma}_{k+1,h}}.$
This update enjoys a computational cost of only $\mathcal{O}(d^{3}U)$, independent of episode count $k$ (Zhang and Sugiyama, 2023; Lee and Oh, 2024). As for storage costs, it avoids the need to store all historical data by updating the feature covariance matrix $\widehat{\Sigma}_{k,h}$ incrementally, requiring only $\mathcal{O}(d^{2})$ storage cost. We note that a dependence on $U$ is introduced in the computational cost. However, as we have discussed in Remark 1, $U$ can be much smaller than the state space size $S$, which is acceptable. This dependency is typical for model-based methods that directly learn transitions, as it need to control the estimation error of transitions, which typically involves a total of $U$ elements. Similar dependencies have been observed in the literature (Hwang and Oh, 2023). Additionally, even for the model-free method (Yang and Wang, 2020; Zhou et al., 2021) which learn value functions, a common assumption is the value $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}V(s^{\prime})$ can be obtained by an Oracle, which depends on $U$ implicitly.
Then, we show the estimator $\widehat{\theta}_{k,h}$ is close to the true parameter $\theta_{h}^{*}$ by the following lemma.
@@
For any $\delta\in(0,1)$, define the confidence set as
$\widehat{\mathcal{C}}_{k,h}=\bigg{\{}\theta\in\Theta\mid\lVert\theta-\widehat{\theta}_{k,h}\rVert_{\widehat{\Sigma}_{k,h}}\leq\sqrt{\lambda}B+\sqrt{\frac{8}{\kappa}d\log\left(1+\frac{kUH}{d\lambda\delta}\right)}\triangleq\widehat{\beta}_{k}\bigg{\}}.$ (5)

### v3

@@ 3 Problem Setup
There exists $10<\kappa\leq\kappa^{*}<$ 1 such that for all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$ and $s^{\prime},s^{\prime\prime}\in\mathcal{S}_{h,s,a}$, it holds that $\inf_{\theta\in\Theta}p_{s,a}^{s^{\prime}}(\theta)p_{s,a}^{s^{\prime\prime}}(\theta)\geq\kappa$ and $p_{s,a}^{s^{\prime}}(\theta_{h}^{*})p_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\geq\kappa^{*}$.
@@
Assumption 1 is similar to the assumption in generalized linear bandit (Filippi et al., 2010) and logistic bandit (Faury et al., 2020; Abeille et al., 2021) to guarantee the Hessian matrix is non-singular.
Finally, we show the claim about the range of the magnitude of $\kappa$ and $\kappa^{*}$.
Claim 1.
It holds that ${1}/{(U\exp(2B))^{2}}\leq\kappa\leq\kappa^{*}\leq{1}/{U^{2}}$.
4 Statistically Efficient Algorithm
The work of Hwang and Oh (2023) first introduced the MNL mixture MDPs and proposed an algorithm with a regret bound of $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$. However, as discussed in Claim 1, it follows that $U^{2}\leq\kappa^{-1}\leq(U\exp(2B))^{2}$, which results in the regret bound scaling polynomially with the number of reachable states $U$. In the worst case, $U$ can be equal to the size of the state space $S$, thereby undermining the motivation for function approximation, which aims to mitigate the dependence on the large state and action spaces. In this section, we address this significant issue by proposing a statistically efficient algorithm that eliminates this dependence in the dominant term of the regret.
4.1 Parameter Estimation
In this section, we first present the parameter estimation method based on the maximum likelihood estimation (MLE) for MNL mixture MDPs. Next, we review the confidence set construction based on the estimated parameters from previous work (Hwang and Oh, 2023). Finally, we propose our new confidence set construction and highlight the improvements it offers over the previous approach.
@@ 4.1 Parameter Estimation
Since the transition parameter $\theta_{h}^{*}$ is unknown, we need to estimate it using the historical data. At episode $k$, we collect a trajectory $\{(s_{k,h},a_{k,h})\}_{h=1}^{H}$, then define the variable: $y_{k,h}\in\{0,1\}^{S_{k,h}}$ where $y_{k,h}^{s^{\prime}}=\mathds{1}_{\{s^{\prime}=s_{k,h+1}\}}$ for $s^{\prime}\in\mathcal{S}_{k,h}\triangleq\mathcal{S}_{s_{k,h},a_{k,h}}$ and $S_{k,h}=\lvert\mathcal{S}_{k,h}\rvert$. We denote by $p_{k,h}^{s^{\prime}}(\theta)=p_{s_{k,h},a_{k,h}}^{s^{\prime}}(\theta)$. Then $y_{k,h}$ is a sample from the following multinomial distribution:
$y_{k,h}\sim\textnormal{multinomial}(1,[p_{k,h}^{s_{1}}(\theta^{*}),\ldots,p_{k,h}^{s_{S_{k,h}}}(\theta^{*})]),$
@@
where the parameter $1$ indicates that $y_{k,h}$ is a single-trial sample. Furthermore, we define the noise $\epsilon_{k,h}^{s^{\prime}}=y_{k,h}^{s^{\prime}}-p_{k,h}^{s^{\prime}}(\theta_{h}^{*})$. It is clear that $\epsilon_{k,h}\in[-1,1]^{S_{k,h}}$, $\mathbb{E}[\epsilon_{k,h}]=\mathbf{0}$ and $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\epsilon_{i,h}^{s^{\prime}}=0$.
We estimate the parameter $\theta_{h}^{*}$ using the MLE and construct the estimator $\widehat{\theta}_{k,h}$ as follows:
@@
$\widehat{\theta}_{k,h}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\mathcal{L}_{k,h}(\theta)\triangleq\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}-y_{i,h}^{s^{\prime}}\log p_{i,h}^{s^{\prime}}(\theta)+\frac{\lambda_{k}}{2}\lVert\theta\rVert_{2}^{2}.$ (1)
where $\lambda_{k}$ is the regularization parameter. Though the MLE estimator $\widehat{\theta}_{k,h}$ is the same as that of Hwang and Oh (2023), the confidence set is constructed differently. Specifically, define the gradient $\mathcal{G}_{k,h}(\theta)$ and Hessian matrix $\mathcal{H}_{k,h}(\theta)$ of the MLE loss by
$\mathcal{G}_{k,h}(\theta)$ $=\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(p_{i,h}^{s^{\prime}}(\theta)-y_{i,h}^{s^{\prime}})\phi_{i,h}^{s^{\prime}}+\lambda_{k}\theta,$
$\mathcal{H}_{k,h}(\theta)$ $=\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\theta)\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime}})^{\top}-\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\theta)p_{i,h}^{s^{\prime\prime}}(\theta)\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime\prime}})^{\top}+\lambda_{k}I.$
Furthermore, we define the feature covariance matrix $A_{k,h}=\kappa^{-1}\lambda_{k}I+\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime}})^{\top}$. By demonstrating $\mathcal{H}_{k,h}(\theta)\succeq\kappa A_{k,h},\forall\theta\in\Theta$, Hwang and Oh (2023) construct the confidence set as
$\mathcal{C}_{k,h}=\Big{\{}\big{\lVert}\theta-\widehat{\theta}_{k,h}\big{\rVert}_{A_{k,h}}\leq\kappa^{-1}\sqrt{d\log(kH/\delta)}\triangleq\beta_{k}\Big{\}}.$ (2)
Since the radius of the confidence set depends on $\kappa^{-1}$, the final regret bound also exhibits a dependence on $\kappa^{-1}$. To eliminates this dependence, we construct a $\kappa$-independent confidence set based on new Bernstein-like inequalities in Lemma 13, following recent advances in logistic bandits (Faury et al., 2020; Périvier and Goyal, 2022). Specifically, we show the following lemma.
@@
For any $\delta\in(0,1)$, set $\lambda_{k}=d\log(kH/\delta)$ and define the confidence set as
$\widehat{\mathcal{C}}_{k,h}=\bigg{\{}\theta\in\Theta\mid\big{\lVert}\mathcal{G}_{k,h}(\theta)-\mathcal{G}_{k,h}(\widehat{\theta}_{k,h})\big{\rVert}_{\mathcal{H}_{k,h}^{-1}(\theta)}\leq(B+3)\sqrt{d\log(kH/\delta)}\triangleq\widehat{\beta}_{k}\bigg{\}}.$ (3)
@@
Comparison to prior work.  We compare the confidence sets defined in Eqs. (2) and (3).
For the confidence set in (3), by the self-concordance property of log-loss in Lemma 10, we have:
$\big{\lVert}\theta-\widehat{\theta}_{k,h}\big{\rVert}_{\mathcal{H}_{k,h}(\theta)}\leq(1+3\sqrt{2})\big{\lVert}\mathcal{G}_{k,h}(\theta)-\mathcal{G}_{k,h}(\widehat{\theta}_{k,h})\big{\rVert}_{\mathcal{H}_{k,h}^{-1}(\theta)}\leq(1+3\sqrt{2})\widehat{\beta}_{k}.$
Then, note that $\mathcal{H}_{k,h}(\theta)\succeq\kappa A_{k,h}$ for all $\theta\in\Theta$, we have
$\big{\lVert}\theta-\widehat{\theta}_{k,h}\big{\rVert}_{A_{k,h}}\leq\kappa^{-1/2}\big{\lVert}\theta-\widehat{\theta}_{k,h}\big{\rVert}_{\mathcal{H}_{k,h}(\theta)}\leq\kappa^{-1/2}(1+3\sqrt{2})(B+3)\sqrt{d\log(kH/\delta)}.$ (4)
Thus, compared to the confidence set in Eq. (2) from Hwang and Oh (2023), our confidence set in Eq. (3) provides a strict improvement by at least a factor of $\kappa^{-1/2}$. This improvement is one of the key components to eliminate the dependence on $\kappa^{-1}$ in the dominant term of the final regret bound.
Additionally, we identify a technical issue of Hwang and Oh (2023). Specifically, they bound the confidence set in Eq. (2) using the self-normalized concentration in Lemma 12. However, the noise is not independent, and since $\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\epsilon_{i,h}^{s^{\prime}}=0$ (due to the learner visiting each stage $h$ exactly once per episode), it does not satisfy the zero-mean sub-Gaussian condition in Lemma 12. We observe similar oversights in multinomial logit contextual bandits (Oh and Iyengar, 2019, 2021; Agrawal et al., 2023), an issue that, to our knowledge, has not been explicitly addressed in prior work. This issue can be resolved with only slight modifications in constant factors by a new self-normalized concentration with dependent noises in Lemma 1 of Li et al. (2024b), a simplified version of Lemma 13.
4.2 Optimistic Value Function Construction
Given the confidence set $\widehat{\mathcal{C}}_{k,h}$, it is natural to follow the principle of “optimism in the face of uncertainty” and construct the optimistic value function. Hwang and Oh (2023) constructed the optimistic value function $\bar{Q}_{k,h}(s,a)$ by adding a closed-form upper confidence bound as follows:
$\bar{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\bar{V}_{k,h+1}(s^{\prime})+2H\beta_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{A_{k,h}^{-1}}\bigg{]}_{[0,H]},$ (5)
where $\bar{V}_{k,h}(s)=\max_{a\in\mathcal{A}}\bar{Q}_{k,h}(s,a)$. Then, a naive idea to compute the optimistic value function is replacing the radius of the confidence set $\beta_{k}$ with $\widehat{\beta}_{k}$ and the feature covariance matrix $A_{k,h}$ with the Hessian matrix $\mathcal{H}_{k,h}(\theta_{h}^{*})$. However, there are two issues with this approach. First, the true parameter $\theta_{h}^{*}$ is unknown thus the Hessian matrix $\mathcal{H}_{k,h}(\theta_{h}^{*})$ is not computable in the algorithmic updates. Second, though $\widehat{\beta}_{k}$ is independent of $\kappa$ and the Hessian matrix $\mathcal{H}_{k,h}(\theta_{h}^{*})$ captures local information, the bonus term $\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}(\theta)}$ remains in a global form. This term involves taking the maximum over all states $s^{\prime}\in\mathcal{S}_{h,s,a}$, which prevents fully utilizing the local information.
To address these challenges, we construct the optimistic value function by directly taking the maximum expected reward over the confidence set. Specifically, we define $\widehat{Q}_{k,h}(s,a)$ and $\widehat{V}_{k,h}(s)$ as
This construction addresses the first challenge by eliminating the need for the Hessian matrix $\mathcal{H}_{k,h}(\theta_{h}^{*})$ and directly leveraging the local information embedded in the confidence set $\widehat{\mathcal{C}}_{k,h}$. For the second challenge, although we bypass this issue in the construction of the optimistic value function, we still need to address it in the analysis. To tackle this, we employ a second-order Taylor expansion, in contrast to the first-order expansion used in the analysis of Hwang and Oh (2023). This allows for a more precise capture of local information. Further details are provided in Lemma 7 in the appendix.
4.3 Regret Guarantee
Algorithm 1 UCRL-MNL-LL
1:  Initialization: Set $\widehat{\theta}_{1,h}=\mathbf{0},\widehat{Q}_{1,h}(\cdot,\cdot)=0,\widehat{V}_{1,h}(\cdot)=0$ for all $h\in[H]$.
2:  for $k=1,\ldots,K$ do
3:     for $h=1,\ldots,H$ do
4:        Observe current state $s_{k,h}$ and select action $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$.
5:     end for
6:     Set $\widehat{V}_{k+1,H+1}(\cdot)=0$.
7:     for $h=H,\ldots,1$ do
8:        Compute the estimator $\widehat{\theta}_{k+1,h}$ by Eq. (1) and update the confidence set $\widehat{\mathcal{C}}_{k+1,h}$ by Eq. (3).
9:        Compute $\widehat{Q}_{k+1,h}(\cdot,\cdot)$ and $\widehat{V}_{k+1,h}(\cdot)$ as in Eq. (6).
10:     end for
11:  end for
Based on the parameter estimation in Section 4.1 and the construction of the optimistic value function in Section 4.2, we propose the UCRL-MNL-LL algorithm. At each stage $h$ of episode $k$, the algorithm observes the current state $s_{k,h}$ and selects the action that maximizes the value function, i.e., $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$, and transits to next state $s_{k,h+1}$. After collecting the trajectory $\{s_{k,h},a_{k,h}\}_{h=1}^{H}$, the estimator $\widehat{\theta}_{k+1,h}$ is updated using Eq. (1), and the confidence set $\widehat{\mathcal{C}}_{k+1,h}$ is updated according to Eq. (3). Then, the value function $\widehat{Q}_{k+1,h}$ and $\widehat{V}_{k+1,h}$ are updated using Eq. (6). The detailed procedure is outlined in Algorithm 1. We show it achieves the following regret guarantee.
Theorem 1.
For any $\delta\in(0,1)$, set $\lambda_{k}=d\log(kH/\delta)$, and $\widehat{\beta}_{k}=(B+3)\sqrt{d\log(kH/\delta)}$. With probability at least $1-\delta$, UCRL-MNL-LL algorithm (Algorithm 1) ensures the following guarantee: